# Decoupled Training: Return of Frustratingly Easy Multi-Domain Learning

### v1

@@ Introduction
In this section, we adopt a popular dataset named Amazon Product Review (Amazon [^3] https://jmcauley.ucsd.edu/data/amazon/. We select $7$ typical subsets with various scales including Books (Books), Electronics (Elec.), Movies_and_TV (TV), CDs_and_Vinyl (CD), Kindle_Store (Kindle), Cell_Phones_and_Accessories (Phone), Digital_Music (Music). As shown in Tabel 4, different domains have various samples from $0.38M$ to $19.2M$. For each domain, diverse user behaviors are available, including more than $5$ reviews for each user-goods pair. The features used for experiments consist of goods_id and user_id. It is obvious that users in these domains have different preferences for various goods. D-Train and all baselines in this section are implemented based on a popular open-sourced library named pytorch-fm  [^4] https://github.com/rixwew/pytorch-fm. We use DNN as the CTR method and the embed_dim and mlp_dim are both set as $16$. The layer number of the expert and the tower are set as $2$ and $3$ respectively. We report AUC (Area Under the Curve) for each domain. Further, AUC_d and AUC_s are averaged over all domains and all samples respectively to intuitively compare D-Train with other baselines. For all models, we use Adam as the optimizer with exponential decay, in which the learning rate starts at $1e^{-3}$ with a decay rate of $1e^{-6}$. During training, the mini-batch size is set to $2048$. As shown in Table 4, D-Train yields larger improvements than a variety of MDL baselines on Amazon Product Review dataset.
@@
D-Train can be used as a general plug-in unit for existing MDL methods. We compare its performance with Shared Bottom, MMoE, and PLE, by only training the parameters of domain-specific heads while fixing the other parameters at the fine-tuning phase. As shown in Table 4, D-Train can further improve these competitive MDL methods on Amazon Product Review dataset, by tailoring D-Train into them.
@@
As shown in Table 2, we conduct an ablation study on DomainNet by removing each phase respectively. The reported results on four different domains reveal that only utilizing all of these phases works best. In particular, D-Train (w/o Fine-tune) works much worse than other ablation experiments, which reveals the importance of decoupling training for domain independence.
@@ Conclusion
Multi-domain learning (MDL) is of great importance in both academia and industry. Numerous efforts have been paid for it but most of the works focus on a special application and a general approach is missing. In this paper, we explicitly uncover the main challenges of dataset bias and domain domination in multi-domain learning, especially the latter since it is usually ignored in most existing works. We further propose a frustratingly easy and hyperparameter-free multi-domain learning method named Decoupled Train (D-Train) with a tri-phase general-to-specific training strategy. Despite its extraordinary simplicity and efficiency, D-Train performs remarkably well in extensive evaluations of various datasets from standard benchmarks to applications of satellite imagery and recommender system.
Appendix A Experiment Details
Baselines
@@ Appendix A Experiment Details
Method #Parameters Hyperparameters
Separately Train $\mathcal{O}(T\cdot n_{\psi}+T\cdot n_{h})$ W/O
@@
MulANN $\mathcal{O}(n_{\psi}+n_{h}+n_{D})$ loss trade-off: $\lambda$; domain discriminator arch.
@@
Shared Bottom $\mathcal{O}(n_{\psi}+T\cdot n_{h})$ W/O
@@
D-Train (ours) $\mathcal{O}(n_{\psi}+T\cdot n_{h})$ W/O
Datasets
@@
Appendix B Parameters and Hyperparameters

### v2

@@ Introduction
We adopt a popular dataset named Amazon Product Review (Amazon) which is a large crawl of product reviews from the Amazon website, recording users’ preferences (by score) for different products. We select $7$ typical subsets with various scales including Books (Books), Electronics (Elec.), Movies_and_TV (TV), CDs_and_Vinyl (CD), Kindle_Store (Kindle), Cell_Phones_and_Accessories (Phone), Digital_Music (Music). As shown in Tabel 4, different domains have various samples from $0.38M$ to $19.2M$. For each domain, diverse user behaviors are available, including more than $5$ reviews for each user-goods pair. The features consist of goods_id and user_id. Users in these domains have different preferences for various goods.
We use DNN as the CTR method and the embed_dim and mlp_dim are both set as $16$. The layer number of the expert and the tower are set as $2$ and $3$ respectively. We report AUC (Area Under the Curve) for each domain. Further, AUC_d and AUC_s are averaged over all domains and all samples respectively to intuitively compare D-Train with other baselines. For all models, we use Adam as the optimizer with exponential decay, in which the learning rate starts at $1e^{-3}$ with a decay rate of $1e^{-6}$. During training, the mini-batch size is set to $2048$. As shown in Table 4, D-Train yields larger improvements than a variety of MDL baselines.
@@
D-Train can be used as a general plug-in unit for existing MDL methods. When D-Train is used as a plugin unit, MMoE and PLE add an extra phase to only train the parameters of domain-specific heads while fixing the other parameters. As shown in Table 4, D-Train can further improve these competitive MDL methods on Amazon Product Review dataset, by tailoring D-Train into them.
@@
As shown in Table 2, ablation study on DomainNet by removing each phase respectively reveals that only utilizing all of these phases works best. In particular, D-Train (w/o Fine-tune) works much worse than other ablation experiments, which reveals the importance of decoupling training for domain independence.
Online Development
We empirically evaluate D-Train in the online advertising scenario of Tencent’s DSP, which needs to serve billions of requests per day. MMoE-like architecture is utilized to predict pCVR for various conversion tasks, such as purchase and follow. Hundreds of features are used in total, including user behavior features, ad side features, and context side features. Online A/B testing from early July 2022 to August 2022 demonstrates that D-Train achieves $0.36\%$ cost lift and $1.69\%$ GMV(normal) lift over the online baseline, where the improvements on different domains are $3.71\%$, $-0.08\%$ (not statistically significant) and $2.03\%$, respectively.
@@ Conclusion
Multi-domain learning (MDL) is of great importance in both academia and industry. In this paper, we explicitly uncover the main challenges of dataset bias and domain domination in multi-domain learning, especially the latter since it is usually ignored in most existing works. We further propose a frustratingly easy and hyperparameter-free multi-domain learning method named Decoupled Train (D-Train) that highlights the domain-independent fine-tuning to alleviate the obstacles of seesaw effect across multiple domains.
Acknowledgements
We would like to thank many colleagues, in particular, Junguang Jiang, Yiwen Qiu, and Prof. Mingsheng Long for their valuable discussions and efforts. Further, we would like to thank Rong Chen, Yuxuan Han, and Wei Xue for their successful deployments of this work into our online system, with a significant improvement in the online A/B test.
Appendix A A. Experiment Details
A.1. Baselines
@@
Type Method #Parameters Hyperparameters
Single Branch Separately Train $\mathcal{O}(T\cdot n_{\psi}+T\cdot n_{h})$ W/O
@@
Domain Alignment MulANN $\mathcal{O}(n_{\psi}+n_{h}+n_{D})$ loss trade-off: $\lambda$; domain discriminator arch.
@@
Multiple Branch Shared Bottom $\mathcal{O}(n_{\psi}+T\cdot n_{h})$ W/O
@@
The Proposed D-Train $\mathcal{O}(n_{\psi}+T\cdot n_{h})$ W/O
A.2. Datasets
@@
Appendix B B. Parameters and Hyperparameters